OpenStudy (j3nny0rtiz):

solve the system of equations by graphing. -1/3x+y=-1 y=4+1/3x

OpenStudy (j3nny0rtiz):

please help

OpenStudy (mathmale):

for clarity, please enclose your fractional coefficient within parentheses: (-1.3(x + y = -1. do the same for the other equation. Next, seeing that both equations have a term with 3 as the denominator, multiply both equations by 3 to eliminate the fractions. Mind showing your work?

OpenStudy (mathmale):

Can you now graph the resulting equations?

OpenStudy (mathmale):

Excuse me; I meant to type (-1/3)x + y = -1. Multiply every term by 3 to clear out the fractions.

OpenStudy (mathmale):

alternatively, you could find the x- and y-intercepts of the line represented by this equation. To find the x-intercept, let y = 0. then (-1/3)x + 0 = -1. What is x? To find the y-intercept, let x = 0. Then 0 + y = -1. What is y? y-intercept is ( ? , ? )

OpenStudy (mathmale):

sorry, Jen, but I have to get off my computer. perhaps we'll find one another both online at the same time and thus be able to continue this discussion. In the meantime, please try to graph the 2 equations on the same set of coordinate axes.